In 4 years you would like to have $5141 for a down payment for a car. The local bank is paying 2% interest, compounded annually. How much would you need to deposit today to reach your goal of $5141? Enter your answer as follows: 1234 Round your answer. Do not use a dollar sign($, a comma( or any decimal points()
Answer #1

Solution: 4749

Working:

Given: FV = $5141; n =4; i = 2%

PV = ?

FV = PV (1 + R/100)^n

5141 = PV ( 1 + 0.02)^4

PV = 5141 / 1.08 = 4749.49 or 4749